OpenStudy (anonymous):

How do I solve 81-4x^2=0

OpenStudy (anonymous):

steps: 1) add 4x^2 to both side 2) divide both sides by 4 3) take square root both sides you will get a value for x that is the solution

OpenStudy (anonymous):

Okay thank you :D, but is it just the one value for x?

OpenStudy (anonymous):

Also this is a non calculator question so how do I square root 20.25?

OpenStudy (anonymous):

in this case x could have two possible values, one positive and one negative.

OpenStudy (anonymous):

write 20.25 like 2025/100

OpenStudy (anonymous):

then find the square root of 2025/100

OpenStudy (uri):

81-4x^2=0 Bring 81 on the right side 4x^2=81 Divide 81 with 4 x^2=81/4

OpenStudy (uri):

x=(+/-) sqrt(81/4) x=(+/-) 9/2

OpenStudy (anonymous):

Okay thank you so much :D

OpenStudy (uri):

Yw,cheese ball.